The attendance at a scholol variety show was 578 people. A ticket for the lower box is worth 75 and a ticket for the upper box is worth 40. If the total revenue is 29,295,about how many tickets were sold for the lower box?




(Ans. algebrically)

Answer by mananth(16946)

lower box - x tickets --------------75 pesos
upper box -578-x tickets------------40 pesos
..
total collection = 29,295
..
75x+40(578-x)=29,295
75x+23,120-40x=29,295
35x+23120=29295
35x=29,295-23120
35x=6175
x=6175/35
x= 176 tickets
...
you can also set up a linear system
x+y = 578
75x+40y = 29295
...
m.ananth@hotmail.ca